# Exploring Pretraining via Active Forgetting for Improving Cross Lingual Transfer for Decoder Language Models

### v1

@@ Abstract
Divyanshu Aggarwalthanks: Joint first authors, Ashutosh Sathefootnotemark: thanks: work done while at Interning at Microsoft, Sunayana Sitaram Microsoft Research India Contact: t-daggarwal@microsoft.com
@@ 2 Related Work
Earlier works have shown that crosslingual transfer can be beneficial for multilingual pretrained models to gain task ability from English-only labelled data Rajaee and Monz (2024); Deb et al. (2023); Parović et al. (2023); Zhao et al. (2024b). While these techniques are effective to a certain extent, they are limited by the multilingual abilities of the pretrained model. Moreover, they cannot be effective for the low resource settings even on multilingual models due to poorer representation of rarer tokens in pretraining corpus and high token fertility 111for morphologically richer languages.
@@ 5 Discussion
We study how language adaptation affects performance on each language class 333(“pretraining”, “adapting” or “other”) by studying performance on translation in Table 3. We find that BA models show significantly better performance (as compared to Baseline) on ‘‘adapting’’ languages at all model scales. Moreover, improvement of BA over Baseline is larger than improvement of AFA over Baseline for larger models and BA models seem to degrade performance on all other language classes leading to worse overall($\mu_{\text{overall}}$). 444 scores as observed in Figure 3.

### v2

@@ Abstract
Divyanshu Aggarwalthanks:  Equal contribution   Ashutosh Sathethanks:  Work done during internship at MSRI11footnotemark: 1   Sunayana Sitaram Microsoft Research India t-daggarwal@microsoft.com, absathe@cse.iitb.ac.in, sunayana.sitaram@microsoft.com
@@ 2 Related Work
Earlier works have shown that crosslingual transfer can be beneficial for multilingual pretrained models to gain task ability from English-only labelled data Rajaee and Monz (2024); Deb et al. (2023); Parović et al. (2023); Zhao et al. (2024b). While these techniques are effective to a certain extent, they are limited by the multilingual abilities of the pretrained model. Moreover, they cannot be effective for the low resource settings even on multilingual models due to poorer representation of rarer tokens in pretraining corpus and high token fertility for morphologically richer languages.
@@ 5 Discussion
We study how language adaptation affects performance on each language class (‘‘pretraining’’, ‘‘adapting’’ or ‘‘other’’) by studying performance on translation in Table 3. We find that BA models show significantly better performance (as compared to Baseline) on ‘‘adapting’’ languages at all model scales. Moreover, improvement of BA over Baseline is larger than improvement of AFA over Baseline for larger models and BA models seem to degrade performance on all other language classes leading to worse overall($\mu_{\text{overall}}$) scores as observed in Figure 3.